Find the area of the surface. The part of the plane z = 6 + 2x + 3y that lies above the rectangle (0, 9] [1, 6]
Answer #1

solution :- Z = 6+ 2x +31 Als) = S( VI + [fa] ² + [fr] 2 dady z = f(a, y) = 6 + 2x + 37 fa 2 fya - 3 (fo)² - 4 ACD = Storg de